# FAIR Universe HiggsML Uncertainty Challenge Competition

### v3

@@ Acknowledgements
We are grateful to the US Department of Energy, Office of High Energy Physics, subprogram on Computational High Energy Physics, for sponsoring this research, as well as to the ANR Chair of Artificial Intelligence HUMANIA (ANR-19-CHIA-0022). Seminal discussions contributing to this work took place at the workshop “Artificial Intelligence and the Uncertainty Challenge in Fundamental Physics,” sponsored by the CNRS AISSAI Center and the DATAIA Institute, and hosted at Institut Pascal at Université Paris-Saclay. The DATAIA Institute and Institut Pascal are respectively funded by the “Investissements d’Avenir” programs ANR-17-CONV-003 and ANR-11-IDEX-0003-01.
@@ Appendix A Proton collisions and detection
This appendix gives a very minimal introduction to what the Challenge data is about (primarily taken from higgsmldoc ).
@@
The different types of particles or pseudo particles of interest for the challenge are electrons, muons, hadronic tau, jets, and missing transverse energy. Electrons, muons, and taus are the three leptons [^5] For the list of elementary particles and their families, we refer the reader to http://www.sciencemag.org/content/338/6114/1558.full. from the standard model.
Electrons and muons live long enough to reach the detector, so their properties (energy and direction) can be measured directly. Conversely, Taus decay almost immediately after their creation into either an electron and two neutrinos, a muon and two neutrinos, or a bunch of hadrons (charged particles) and a neutrino. The bunch of hadrons can be identified as a pseudo particle called the hadronic tau. Jets are pseudo particles rather than real particles; they originate from a high-energy quark or gluon and appear in the detector as a collimated energy deposit associated with charged tracks. The measured momenta (see Appendix B for a short introduction to special relativity) of all the particles of the event is the primary information provided for the challenge.
@@
To summarize, for each event, we produce a list of momenta for an electron or muon, a tau hadron, up to two jets, plus the missing transverse energy.
@@ Appendix B Special relativity
This appendix gives a very minimal introduction to special relativity for a better understanding of how the Higgs boson search is performed and what the extracted features mean (taken mainly from higgsmldoc ).
@@ B.1 Momentum, mass, and energy
The relation holds both for the norms $v$ and $p$ and for the three dimensional vectors $\vec{v}$ and $\vec{p}$, that is, $\vec{p}=\gamma m\vec{v}$, where, by convention, $p=|\vec{p}|$ and $v=|\vec{v}|$. The factor $\gamma$ diverges to infinity when $v$ is close to $c$, and the speed of light cannot be reached nor surpassed. Hence, momentum is a concept more frequently used than speed in particle physics. The kinematics of a particle is fully defined by the momentum and energy, more precisely, by the 4-momentum $(p_{x},p_{y},p_{z},E)$. When a particle is identified, it has a well-defined mass666neglecting the particle width, so its energy can be computed from the momentum and mass using Equation 8. Conversely, the mass of a particle with known momentum and energy can be obtained from
@@ B.2 Invariant mass
The mass of a particle is an intrinsic property of a particle. So, for all events with a Higgs boson, the Higgs boson will have the same mass. To measure the mass of the Higgs boson, we need the 4-momentum $(p_{x},p_{y},p_{z},E)=(\vec{p},E)$ of its decay products. Take the simple case of the Higgs boson $H$ decaying into a final state of two particles, $A$ and $B$, which are measured in the detector. By conservation of the energy and momentum (which are fundamental laws of nature), we can write $E_{H}=E_{A}+E_{B}$ and $\vec{p}_{H}=\vec{p}_{A}+\vec{p}_{B}$. Since the energies and momenta of $A$ and $B$ are measured in the detector, we can compute $E_{H}$ and $p_{H}=|\vec{p}_{H}|$ and calculate $m_{H}=\sqrt{E_{H}^{2}-p_{H}^{2}}$. This is called the invariant mass because (with a perfect detector) $m_{H}$ remains the same even if $E_{H}$ and $p_{H}$ differ from event to event. This can be generalized to more than two particles in the final state and to any number of intermediate states.
In our case, the final state for particles originating from the Higgs boson is a lepton, a hadronic tau, and three neutrinos. The lepton and hadronic tau are measured in the detector, but for the neutrinos, all we have is the transverse missing energy, which estimates the sum of the momenta of the three neutrinos in the transverse plane. Hence, the mass of the $\tau\tau$ can not be measured; we have to resort to different estimators which are only correlated to the mass of the $\tau\tau$. For example, the visible mass (feature DER_mass_vis) which is the invariant mass of the lepton and the hadronic tau, hence deliberately ignoring the unmeasured neutrinos. The possible jets in the events are not originating from the Higgs boson itself but can be produced in association with it.
@@ Appendix C The detailed description of the features
Prefix-less variables Weight, Label,DetailedLabel, have a special role and should not be used as regular features for the model [^7] In the starting kit they are split away in separated numpy arrays while the regular features are stored in a Dataframe:
@@
The variables prefixed with PRI (for PRImitives) are “raw” quantities about the bunch collision as measured by the detector, essentially parameters of the momenta of particles.
@@
Variables prefixed with DER (for DERived) are quantities computed from the primitive features on the fly from PRImary features (including possible systematics shifts ) [^8] The code to compute DERived features from PRImitive features can be seen at https://github.com/FAIR-Universe/HEP-Challenge/blob/master/ingestion_program/derived_quantities.py. These quantities were selected by the physicists of ATLAS in the reference document AtlasHiggsInTauTauPaper15 either to select regions of interest or as features for the Boosted Decision Trees used in this analysis in order to enhance signal Higgs boson events separation from background events. DERived features were already present in the HiggsML dataset higgsmldoc  [^9] The notable exception of DER_mass_MMC which was in the HiggsML dataset but is deliberately absent from the Fair-Universe dataset because it was the result of a complex and lengthy Monte-Carlo Markov Chain integration which is not practical to rerun.). The DERived features correspond to feature engineering; an ideal model to be trained on infinite statistics should not need these features. This distinction between primary and derived features (or "low-level" and "high-level" or "raw variables" and "human-assisted variables") is rather standard in the AI for HEP literature, see for example BaldiExotic ; BaldiTauTau . There is no guarantee that all DERived features are useful for this challenge (they could even be detrimental in the context of systematics). The challenge participant is free to keep these DERived features, remove them altogether, keep a few, or do more feature engineering.
@@
The feature list and event sample are primarily inspired from AtlasHiggsInTauTauPaper15 . One crucial difference is that the dataset was produced with a more straightforward (leading-order) event generator (Pythia), and the detector effect was simulated with a more straightforward detector simulation (Delphes rather than Geant4 ATLAS Simulation). These simplifications allowed us to provide to participants a large sample allowing the development of sophisticated models while preserving the complexity of the original problem.
@@ Appendix D Systematic biases
This appendix details the implementation of the systematic biases Nuisance Parameters [^10] See also https://github.com/FAIR-Universe/HEP-Challenge/blob/master/ingestion_program/systematics.py.
@@ D.1 Systematic bias definition
Table 2 lists the different Nuisance Parameter with their Gaussian distribution and the range to which they are clipped. $\alpha_{\rm tes}$, $\alpha_{\rm jes}$, and $\alpha_{\rm soft\_met}$ impacts some PRImary features, and then DERived features in cascade. $\alpha_{\rm tes}$ and $\alpha_{\rm jes}$ also impact which events make it to the final dataset. $\alpha_{\rm ttbar\_scale}$, $\alpha_{\rm diboson\_scale}$ and $\alpha_{\rm bkg\_scale}$ only impact the Weight of some background categories, that is to say, the composition of the background (for $\alpha_{\rm ttbar\_scale}$ and $\alpha_{\rm diboson\_scale}$) or the overall level of the background $\alpha_{\rm bkg\_scale}$. The Gaussian distributions parameterise our ignorance of the exact value of the biases. We think their value is 1 (or zero for $\alpha_{\rm soft\_met}$) while their real value is slightly different, as parameterised by their width, thus biasing our measurement by an unknown amount, which can be simulated.
@@
Table 2: List of six systematic bias Nuisance Parameter defined in the challenge, with the mean and sigma of their Gaussian (Log-normal for $\alpha_{\rm soft\_met}$) distribution and their range. The corresponding $\alpha$ is set to the Mean value whenever a systematic bias is switched off. "No systematics" means all $\alpha$ are set to their Mean value.
@@ D.4 Event selection
To alleviate this, "high thresholds" (see Table 3) have been defined, which should systematically be applied after the calculation of the biased PRImary parameters, so that the thresholds to be observed on PRI_had_pt, PRI_jet_leading_pt PRI_jet_subleading_pt are independent of $\alpha_{\rm tes}$ and $\alpha_{\rm jes}$. The ranges in Table 2 are such that the thresholds should also be applied when no systematics bias is used [^11] In practice, function systematics in https://github.com/FAIR-Universe/HEP-Challenge/blob/master/ingestion_program/systematics.py should always be used, even in the no systematics case..

### v4

@@ Acknowledgements
We are grateful to the US Department of Energy, Office of High Energy Physics, and the subprogram on Computational High Energy Physics, for sponsoring this research, as well as to the ANR Chair of Artificial Intelligence HUMANIA (ANR-19-CHIA-0022). Seminal discussions contributing to this work took place at the workshop “Artificial Intelligence and the Uncertainty Challenge in Fundamental Physics,” sponsored by the CNRS AISSAI Center and the DATAIA Institute, and hosted at Institut Pascal at Université Paris-Saclay. The DATAIA Institute and Institut Pascal are respectively funded by the “Investissements d’Avenir” programs ANR-17-CONV-003 and ANR-11-IDEX-0003-01. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award HEP-ERCAP0032917. The computational results of subsection 5.1 [76] were obtained using the CLIP cluster.
@@ Appendix A Proton collisions and detection
This appendix gives details on how the data was generated.
@@
An online trigger system discards most of the bunch collisions containing uninteresting events. The trigger is a three-stage cascade classifier which decreases the event rate from 40 000 00040\,000\,000 to about $400$ per second. The selected $400$ events are saved on disk, producing about one billion events and three petabytes of raw data per year.
The different types of particles or pseudo-particles of interest for the challenge are electrons, muons, hadronic tau, jets, and missing transverse energy. Electrons, muons, and taus are the three leptons [^2] For the list of elementary particles and their families, we refer the reader to http://www.sciencemag.org/content/338/6114/1558.full. from the standard model.
Electrons and muons live long enough to reach the detector, so their properties (energy and direction) can be measured directly. Conversely, Taus decay almost immediately after their creation into either an electron and two neutrinos, a muon and two neutrinos, or a bunch of hadrons (charged particles) and a neutrino. The bunch of hadrons can be identified as a pseudo-particle called the hadronic tau. Jets are pseudo particles rather than real particles; they originate from a high-energy quark or gluon and appear in the detector as a collimated energy deposit associated with charged tracks. The primary information provided for the challenge is the measured momenta (see Appendix B for a short introduction to special relativity) of all the particles of the event.
@@
Figure 7: Diagram of the particles in the final state chosen: one lepton, one tau hadron, up to two jets, and the missing transverse momentum vector, see text for details.
To summarise, for each event, we produce a list of momenta for an electron or muon, a tau hadron, up to two jets, plus the missing transverse energy.
Table 1: Summary of the dataset for each category and subcategory. "Number Generated" is the number of events available in the dataset. In contrast, "LHC events" is the average number in this category in a pseudo-experiment corresponding to running of the Large Hadron Collider for 10 fb-1, corresponding to approximately 800 billion inelastic proton collisions, or 2 weeks in summer 2024 conditions.
Process Number Generated LHC Events Label
Higgs 52 040 227 1 015 signal
Z Boson 160 383 358 1 002 395 background
Di-Boson 605 118 3 783 background
$t\bar{t}$ 7 070 398 44 192 background
Table 1 details the number of events of each category in the dataset.
@@ Appendix B Special relativity
This appendix gives a very minimal introduction to special relativity for a a better understanding of how the Higgs boson search is performed and what the extracted features mean (taken mainly from [82]).
@@ B.1 Momentum, mass, and energy
The relation holds both for the norms $v$ and $p$ and for the three dimensional vectors $\vec{v}$ and $\vec{p}$, that is, $\vec{p}=\gamma m\vec{v}$, where, by convention, $p=|\vec{p}|$ and $v=|\vec{v}|$. The factor $\gamma$ diverges to infinity when $v$ is close to $c$, and the speed of light cannot be reached or surpassed. Hence, momentum is a concept more frequently used than speed in particle physics. The kinematics of a particle is fully defined by the momentum and energy, more precisely, by the 4-momentum $(p_{x},p_{y},p_{z},E)$. When a particle is identified, it has a well-defined mass333neglecting the particle width, so its energy can be computed from the momentum and mass using Equation 6. Conversely, the mass of a particle with known momentum and energy can be obtained from
@@ B.2 Invariant mass
The mass of a particle is an intrinsic property of a particle. So, for all events with a Higgs boson, the Higgs boson will have the same mass. To measure the mass of the Higgs boson, we need the 4-momentum $(p_{x},p_{y},p_{z},E)=(\vec{p},E)$ of its decay products. Take the simple case of the Higgs boson $H$ decaying into a final state of two particles, $A$ and $B$, which are measured in the detector. By conservation of energy and momentum (which are fundamental laws of nature), we can write $E_{H}=E_{A}+E_{B}$ and $\vec{p}_{H}=\vec{p}_{A}+\vec{p}_{B}$. Since the energies and momenta of $A$ and $B$ are measured in the detector, we can compute $E_{H}$ and $p_{H}=|\vec{p}_{H}|$ and calculate $m_{H}=\sqrt{E_{H}^{2}-p_{H}^{2}}$. This is called the invariant mass because (with a perfect detector) $m_{H}$ remains the same even if $E_{H}$ and $p_{H}$ differ from event to event. This can be generalised to more than two particles in the final state and to any number of intermediate states.
In our case, the final state for particles originating from the Higgs boson is a lepton, a hadronic tau, and three neutrinos. The lepton and hadronic tau are measured in the detector, but for the neutrinos, all we have is the transverse missing energy, which estimates the sum of the momenta of the three neutrinos in the transverse plane. Hence, the mass of the $\tau\tau$ can not be measured; we have to resort to different estimators which are only correlated to the mass of the $\tau\tau$. For example, the visible mass (feature DER_mass_vis) which is the invariant mass of the lepton and the hadronic tau, hence deliberately ignoring the unmeasured neutrinos. The possible jets in the events are not originating from the Higgs boson itself, but can be produced in association with it.
@@ Appendix C The detailed description of the features
Prefix-less variables Weight, Label,DetailedLabel, have a special role and should not be used as regular features for the model [^4] In the starting kit, they are split away in separate numpy arrays while the regular features are stored in a Dataframe:
@@
The variables prefixed with PRI (for PRImitives) are “raw” quantities about the bunch collision as measured by the detector, essentially parameters of the momenta of particles (see Figure 8, Figure 9 and Figure 10 for their distributions).
@@
Figure 8: Distributions of: (a) hadron $\eta$, (b) hadron $\phi$, (c) hadron $p_{T}$, (d) all jets $p_{T}$, (e) leading jet $\eta$, and (f) leading jet $\phi$. For jet quantities, the left most bin is the default value in the absence of jets.
Figure 9: Distributions of: (a) leading jet $p_{T}$, (b) subleading jet $\eta$, (c) subleading jet $\phi$, (d) subleading jet $p_{T}$, (e) lepton $\eta$, and (f) lepton $\phi$. For jet quantities, the left most bin is the default value in no jet, or only one jet.
Figure 10: Distributions of: (a) lepton $p_{T}$, (b) MET, (c) MET $\phi$, and (d) number of jets.
Variables prefixed with DER (for DERived) are quantities computed from the primitive features on the fly from PRImary features (including possible systematics shifts ) [^5] The code to compute DERived features from PRImitive features can be seen at https://github.com/FAIR-Universe/FAIR_Universe_dataset/blob/main/hep_challenge/derived_quantities.py(see Figure 11 and Figure 12 for their distributions). These quantities were selected by the physicists of ATLAS in the reference document [83] either to select regions of interest or as features for the Boosted Decision Trees used in this analysis in order to enhance signal Higgs boson events separation from background events. DERived features were already present in the HiggsML dataset [82] [^6] The notable exception of DER_mass_MMC which was in the HiggsML dataset but is deliberately absent from the Fair-Universe dataset because it was the result of a complex and lengthy Monte-Carlo Markov Chain integration which is not practical to rerun.). The DERived features correspond to feature engineering; an ideal model to be trained on infinite statistics should not need these features. This distinction between primary and derived features (or "low-level" and "high-level" or "raw variables" and "human-assisted variables") is rather standard in the AI for HEP literature, see for example [84, 85]. There is no guarantee that all DERived features are useful for this challenge (they could even be detrimental in the context of systematics). The challenge participant is free to keep these DERived features, remove them altogether, keep a few, or do more feature engineering.
@@
Figure 11: Distributions of kinematic variables: (a) $\Delta\eta(jet\text{-}jet)$, (b) $\Delta R(had\text{-}lep)$, (c) $lep\ \eta$ centrality, (d) $m(jet\text{-}jet)$, (e) $m_{T}(\text{MET}\text{-}lep)$, and (f) visible mass.
Figure 12: Distributions of: (a) MET $\phi$ centrality, (b) ${\rm prod\ }\eta(jet\text{-}jet)$, (c) $p_{T}^{h}$, (d) $p_{T}(lep\text{/}had)$ ratio, (e) $p_{T}^{\text{tot}}$, and (f) $\sum p_{T}$.
The feature list and event sample are primarily inspired from [83]. One crucial difference is that the dataset was produced with a more straightforward (leading-order) event generator (Pythia), and the detector effect was simulated with a more straightforward detector simulation (Delphes rather than Geant4 ATLAS Simulation). These simplifications allowed us to provide to participants a large sample allowing the development of sophisticated models while preserving the complexity of the original problem.
@@ Appendix D Systematic biases
This appendix details the implementation of the systematic biases Nuisance Parameters [^7] See also https://github.com/FAIR-Universe/FAIR_Universe_dataset/blob/main/hep_challenge/systematics.py.
@@ D.1 Systematic bias definition
Table 2 lists the different Nuisance Parameters with their Gaussian distribution and the range to which they are clipped. $\alpha_{\rm tes}$, $\alpha_{\rm jes}$, and $\alpha_{\rm soft\_met}$ impacts some PRImary features, and then DERived features in cascade. $\alpha_{\rm tes}$ and $\alpha_{\rm jes}$ also impact which events make it to the final dataset. $\alpha_{\rm ttbar\_scale}$, $\alpha_{\rm diboson\_scale}$ and $\alpha_{\rm bkg\_scale}$ only impact the Weight of some background categories, that is to say, the composition of the background (for $\alpha_{\rm ttbar\_scale}$ and $\alpha_{\rm diboson\_scale}$) or the overall level of the background $\alpha_{\rm bkg\_scale}$. The Gaussian distributions parameterise our ignorance of the exact value of the biases. We think their value is 1 (or zero for $\alpha_{\rm soft\_met}$) while their real value is slightly different, as parameterised by their width, thus biasing our measurement by an unknown amount, which can be simulated.
@@
Table 2: List of six systematic bias Nuisance Parameters defined in the challenge, with the mean and sigma of their Gaussian (Log-normal for $\alpha_{\rm soft\_met}$) distribution and their range. The corresponding $\alpha$ is set to the Mean value whenever a systematic bias is switched off. "No systematics" means all $\alpha$ are set to their Mean value.
@@ D.4 Event selection
To alleviate this, "high thresholds" (see Table 3) have been defined, which should systematically be applied after the calculation of the biased PRImary parameters, so that the thresholds to be observed on PRI_had_pt, PRI_jet_leading_pt PRI_jet_subleading_pt are independent of $\alpha_{\rm tes}$ and $\alpha_{\rm jes}$. The ranges in Table 2 are such that the thresholds should also be applied when no systematics bias is used [^8] In practice, function systematics in https://github.com/FAIR-Universe/FAIR_Universe_dataset/blob/main/hep_challenge/systematics.py should always be used, even in the no systematics case..